Can someone please simplfy these 2 equations for me.
dy/dx = (cosx(3+cos2x)+(-2sin2x)(sinx))/(3+cos2x)^2
and
75/(x+12) + 9/x = 3.
Thank you.

Answer by venugopalramana(3286)

Can someone please simplfy these 2 equations for me.
dy/dx = (cosx(3+cos2x)+(-2sin2x)(sinx))/(3+cos2x)^2
IS Y= (cosx(3+cos2x)+(-2sin2x)(sinx))/(3+cos2x)^2 AND YOU WANT TO FIND DIFFERENTIATE TO FIND DY/DX OR REALLY
dy/dx = (cosx(3+cos2x)+(-2sin2x)(sinx))/(3+cos2x)^2
AND YOU WANT TO INTEGRATE TO FIND Y?
PLEASE COPY PROBLEM PROPERLY AND RESUBMIT